Franklin bikes from A to B at 25km/h.
He returns by car at 75km/h.
The bike trip is 2 hours longer than the car trip.
How far is it from A to B?

Answer by ankor@dixie-net.com(12699)

Franklin bikes from A to B at 25km/h.
He returns by car at 75km/h.
The bike trip is 2 hours longer than the car trip.
How far is it from A to B?
:
Let t = car time
then
(t+2) = bike time
:
The two trips are the same distance; write a dist equation: dist = speed * time
Car dist = Bike dist
75t = 25(t+2)
75t = 25t + 50
75t - 25t = 50
50t = 50
t = 1 hr is the car time, then 3 hrs is the bike time, obviously.
:
Find the dist:
75* 1 = 75 km
or
3*25 = 75 km